It takes 60 seconds to walk up a broken escalator and 40 seconds to ride up the same escalator when it is moving.  How long will it take to walk up the same escalator when it is moving?
Answer by ankor@dixie-net.com(22740)

It takes 60 seconds to walk up a broken escalator and 40 seconds to ride up the same escalator when it is moving.
How long will it take to walk up the same escalator when it is moving?
:
Assume the length of the escalator is 60 ft
then
Walking: 1 ft/sec
moving: 60/40 = 1.5 ft/sec
walking & moving 2.5 ft/sec
:
Find the time at 2.5 ft/sec
t = 60/2.5 = 24 seconds walking & moving